OpenStudy (anonymous):

The expression \[ a^2 b^{-3} \over a^{-4} b^2 \] is equivalent to ?

OpenStudy (turingtest):

have you tried to apply \[\large x^{-a}=\frac1{x^a}\]to get all positive exponents?

OpenStudy (anonymous):

so write it as \[ a^2b^{1\3} \over a^{1\4} b^2\]

OpenStudy (turingtest):

not quite... the whole thing flips, not just the exponent\[\large{{a^2\cdot\frac1{b^3}}\over{\frac1{a^4}\cdot b^2}}\]

OpenStudy (turingtest):

which can simplified...

OpenStudy (turingtest):

\[\large{{a^2\cdot\frac1{b^3}}\over{\frac1{a^4}\cdot b^2}}=\frac{a^2\cdot a^4}{b^3\cdot b^2}\]

OpenStudy (anonymous):

so I add the exponent?

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

so my outcome would be \[a^6 \over b^5\]

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

good job :)

OpenStudy (anonymous):

thank you